3. Data on the mileage of 10 randomly selected cars is listed below. The values are ordered for convenience. 20 20 22 23 24 26 26 27 27 39 Find the two-number summary using your calculator. (Mean, Sample standard deviation) [4 pts.] a. (25.4, 5.4 b. Find the five-number summary. [5 pts.] Mcan-25 uxper-limita lbwer-limit-39 lower IAR-2a Are there any outliers based on 1.5 x IQR rule? [5 pts.] The upper limit:-一ND,no numbe、、 S40.5 The lower limit: NDno number is33 Justify your answer!
Answer #1

a)

(X-x)2 29.1600 29.1600 11.5600 5.7600 1.9600 0.3600 0.3600 2.5600 2.5600 184.9600 268.4000 29.822 5.461 (X-X) 5.400 5.400 3.400 2.400 1.400 S.No 20.000 20.000 22.000 23.000 24.000 26.000 26.000 27.000 27.000 39.000 254 25.400 13.600 s-E(x-X),(n-1)= 试试试试 3 1234567890 x n

frm above mean ; std deviation =(25.4 ; 5.46)

b)

minimum =20

first quartile Q1 =np th value =10*0.25 th value =2.5 th value ~ 22

second quartile Q2 =10*0.5 th value =5 th value =(24+26)/2 =25

third quartile Q3 =10*0.75 th value =7.5 th value =27

maximum=39

c)

upper limit =Q3+1.5*(Q3-Q1)=27+1.5*(27-22)=34.5 ; hence 39 is an upper outlier

lower limit -Q1-1.5*(Q3-Q1)=22-1.5*(27-22)=14.5 ;there is no lower outlier